# Mask-Enhanced Autoregressive Prediction: Pay Less Attention to Learn More

### v1

@@ Abstract
Large Language Models (LLMs) are discovered to suffer from accurately retrieving key information. To address this, we propose Mask-Enhanced Autoregressive Prediction (MEAP), a simple yet effective training paradigm that seamlessly integrates Masked Language Modeling (MLM) into Next-Token Prediction (NTP) to enhance the latter’s in-context retrieval capabilities. Specifically, MEAP first randomly masks a small fraction of input tokens and then directly performs the standard next-token prediction autoregressive using a decoder-only Transformer. MEAP eliminates the need for bidirectional attention or encoder-decoder architectures for MLM, incurring no additional computational overhead during pre-training or inference. Intensive experiments demonstrate that MEAP substantially outperforms NTP on key information retrieval and long-context reasoning tasks, while performing on par or better on commonsense reasoning tasks. The benefits of MEAP also extend to supervised fine-tuning, where it shows remarkable advantages in lost-in-the-middle scenarios, outperforming NTP by 11.77% percentage points. Our analysis indicates that MEAP’s effectiveness arises from its ability to promote more distinguishable attention scores by concentrating on a reduced set of non-masked tokens. This mechanism improves the model’s focus on task-relevant signals while mitigating the influence of peripheral context. These findings position MEAP as a promising training paradigm for large language models.Code is provided in the link.
@@ 1 Introduction
Next-token prediction (NTP) (Radford, 2018) is the foundational training objective for many large language models (LLMs), including OpenAI’s GPT series (Brown, 2020). NTP trains models to predict the next word (or token) in a sequence, given all preceding tokens. Its scaling efficiency and exceptional performance in text generation have established it as the dominant paradigm for state-of-the-art LLMs such as GPT-4 (Achiam et al., 2023), LLaMa3 (Dubey et al., 2024), Gemini 1.5 Pro (Team et al., 2024), and DeepSeek-V3 (Liu et al., 2024a). However, recent studies highlight the limitations of NTP-based LLMs in accurately retrieving key information from context (Liu et al., 2024b; Kamradt, 2023).
@@ 3 Mask-Enhanced Autoregressive Prediction
LLM fine-tuning. MEAP can also be extended to fine-tuning scenarios. In this approach, we duplicate the training samples and apply the same random masking strategy to the copied sequences during fine-tuning. The original sequences and their masked counterparts are then combined into a single input sequence to feed into the model. Formally, the next-token prediction with causal attention mechanism models the following objective:
$p_{\theta}(X^{\prime})=\prod_{t=1}^{T}p_{\theta}(x_{t}\mid x_{1},\dots,x_{t-1};x_{1},[mask],\dots,x_{t-1})$
This design addresses a critical concern: input sequences in supervised fine-tuning often contain key information essential for downstream tasks. Directly masking the original sequence risks removing crucial information, potentially compromising the model’s performance on the target tasks. Masking the duplicated sequence incorporates MLM to NTP while avoiding this concern. We choose $P=10\%$ for fine-tuning in our experiment. We only perform MEAP for the QA pair whose answer length exceeds 50, otherwise, we perform the standard NTP for the pair.
@@ 4.1 Pre-training Evaluation
Setup. We follow the Llama architecture (Touvron et al., 2023) as our base model. Specifically, we train 1.1B decoder-only Transformers (Vaswani, 2017) following the setting of Zhang et al. (2024). Our model has 24 layers with 32 attention heads, a hidden size of 2,048, an intermediate hidden size of 5,632, 32 attention heads, and a context length of 4096. We follow the common configurations of LLM components, e.g., Rotary Positional Embedding (RoPE) (Su et al., 2024), Pre-Norm (Ba, 2016) with RMSNorm (Zhang & Sennrich, 2019), SwiGLU (Shazeer, 2020), and Grouped-query Attention (Ainslie et al., 2023). To assess the scalability of MEAP, we increase the training token size from 40B to 60B, and further to 200B.
@@ 4.1.2 Needle-in-a-Haystack Retrieval
Figure 3: Training dynamics comparison between standard pretraining and MEAP framework. Scores are computed using ROUGE-1, measuring unigram overlap between model responses and expected answers.
@@
To demonstrate MEAP’s scalability, we increase the training token size to 40B, 80B, and 200B, reporting the results of needle retrieval in Table 2. The results show that MEAP consistently outperforms NTP across different training scales. At 40B tokens, MEAP achieves 80.2% accuracy, significantly surpassing the baseline’s 65.9%. The performance gap peaks at 60B tokens, with MEAP maintaining steady improvement and reaching 85.8% accuracy. At 200B tokens, MEAP approaches optimal performance, attaining 98.2% accuracy, while the NTP baseline still falls short of 90% accuracy. It is noteworthy that MEAP achieves 85.8% accuracy using just 60B training tokens, whereas NTP requires approximately three times as many (200B tokens) to reach a similar level. This demonstrates MEAP’s superior data efficiency over NTP in key information retrieval.
@@ 4.1.5 Contextual Hallucination Evaluation
Table 4: Accuracy (i.e., free of hallucinations) on text summarization datasets.
@@
NTP 0.09 0.17 0.24
MEAP 0.13 0.19 0.33
Since MEAP improves more accurate key information retrieval, we expect it to suffer less from contextual hallucination. To verify, we evaluate MEAP in reducing contextual hallucinations on three summarization datasets: XSum (Narayan et al., 2018), WikiSum (Cohen et al., 2021), and MultiNews (Fabbri et al., 2019), following Ye et al. (2024). For this setting, we fine-tune the pre-trained models with Alpaca and evaluate them. We compare model-generated and reference summaries using Deepseek-V3 (Liu et al., 2024a) as the hallucination detector across 100 random samples per dataset. As shown in Table 4, our masking strategy achieves a consistent reduction in hallucination rates across all datasets.
@@ 5 Why Does MEAP Work?
While effective, attention mechanisms in LLMs often struggle with long-context understanding, where redundant and non-informative attention is assigned to tokens (Liu et al., 2024b; Li et al., 2024b). A plausible explanation is that the attention module relies on the Softmax function to normalize attention scores within (0, 1), which tends to minimize differences among tokens, especially when training on sequences of thousands of tokens. Furthermore, LLMs exhibit a phenomenon known as attention sinks, where the initial few tokens receive disproportionately high attention scores compared to the rest (Xiao et al., 2023). Collectively, these factors lead to small and nearly indistinguishable attention scores across tokens, which is generally undesirable. When LLMs fail to properly differentiate between tokens, they are more likely to generate incorrect outputs.
@@ 5.2 MEAP Focus More on Task-Relevant Tokens
Table 8: Attention score comparison between NTP and MEAP.
Input Length Mask Ratio Score Decay Var. Increase
1,024 0.15 34.08% 12.66%
4,096 0.15 53.34% 7.80%

### v2

@@ Abstract
Large Language Models (LLMs) are discovered to suffer from accurately retrieving key information. To address this, we propose Mask-Enhanced Autoregressive Prediction (MEAP), a simple yet effective training paradigm that seamlessly integrates Masked Language Modeling (MLM) into Next-Token Prediction (NTP) to enhance the latter’s in-context retrieval capabilities. Specifically, MEAP first randomly masks a small fraction of input tokens and then directly performs the standard next-token prediction autoregressive using a decoder-only Transformer. MEAP eliminates the need for bidirectional attention or encoder-decoder architectures for MLM, incurring no additional computational overhead during pre-training or inference. Intensive experiments demonstrate that MEAP substantially outperforms NTP on key information retrieval and long-context reasoning tasks, while performing on par or better on commonsense reasoning tasks. The benefits of MEAP also extend to supervised fine-tuning, where it shows remarkable advantages in lost-in-the-middle scenarios, outperforming NTP by 11.77% percentage points. Our analysis indicates that MEAP’s effectiveness arises from its ability to promote more distinguishable attention scores by concentrating on a reduced set of non-masked tokens. This mechanism improves the model’s focus on task-relevant signals while mitigating the influence of peripheral context. These findings position MEAP as a promising training paradigm for large language models. Code is available at https://github.com/scitix/MEAP.
@@ 1 Introduction
Next-token prediction (NTP) (Radford, 2018) is the foundational training objective for many large language models (LLMs), including OpenAI’s GPT series (Brown, 2020). NTP trains models to predict the next word (or token) in a sequence, given all preceding tokens. Its scaling efficiency and exceptional performance in text generation have established it as the dominant paradigm for state-of-the-art LLMs such as GPT-4 (Achiam et al., 2023), LLaMa3 (Dubey et al., 2024), Gemini 1.5 Pro (Team et al., 2024), and DeepSeek-V3 (Liu et al., 2024a). However, recent studies highlight the limitations of NTP-based LLMs in accurately retrieving key information from context (Liu et al., 2024b; Kamradt, 2023; Nelson et al., 2024).
@@ 3 Mask-Enhanced Autoregressive Prediction
LLM fine-tuning. MEAP can also be extended to fine-tuning scenarios. In this scenario, we duplicate the training samples and apply the same random masking strategy to the copied sequences during fine-tuning. The original sequences and their masked counterparts are then combined into a single input sequence to feed into the model. The cross-entropy loss is computed only with the masked tokens ($U_{m}$) in the answer tokens ($U_{q}$). This design addresses a critical concern: input sequences in supervised fine-tuning often contain key information essential for downstream tasks. Directly masking the original sequence risks removing crucial information, potentially compromising the model’s performance on the target tasks. Masking the duplicated sequence incorporates MLM to NTP while avoiding this concern. We choose $P=10\%$ for fine-tuning in our experiment. We only perform MEAP for the QA pair whose answer length exceeds 50, otherwise, we perform the standard NTP for the pair. Formally, the following objective of MEAP for fine-tuning is:
$\mathcal{L}(\theta)=-\sum_{t\in U_{q}\cup U_{m}}\log p_{\theta}(x_{t}\mid x_{1},\dots,x_{t-1};\hat{x}_{1},[\text{mask}],\dots,\hat{x}_{t-1})$
where the sequence $\{\hat{x}_{i}\}$ is a copy of the original sequence $\{x_{i}\}$ (i.e., $\hat{x}_{i}=x_{i}$).
@@ 4.1 Pre-training Evaluation
Setup. We follow the Llama architecture (Touvron et al., 2023) as our base model. Specifically, we train 1.1B decoder-only Transformers (Vaswani, 2017) following the setting of Zhang et al. (2024). Our model has 24 layers with 32 attention heads, a hidden size of 2,048, an intermediate hidden size of 5,632, and a context length of 4096. We follow the common configurations of LLM components, e.g., Rotary Positional Embedding (RoPE) (Su et al., 2024), Pre-Norm (Ba, 2016) with RMSNorm (Zhang & Sennrich, 2019), SwiGLU (Shazeer, 2020), and Grouped-query Attention (Ainslie et al., 2023). To assess the scalability of MEAP, we increase the training token size from 40B to 60B, and further to 200B.
@@ 4.1.2 Needle-in-a-Haystack Retrieval
Figure 3: Performance comparison between NTP and MEAP on Needle In A Haystack. Scores are computed using ROUGE-1, measuring unigram overlap between model responses and expected answers.
@@
To demonstrate MEAP’s scalability, we increase the training token size to 40B, 60B, and 200B, reporting the results of needle retrieval in Table 2. The results show that MEAP consistently outperforms NTP across different training scales. At 40B tokens, MEAP achieves 80.2% accuracy, significantly surpassing the baseline’s 65.9%. The performance gap peaks at 60B tokens, with MEAP maintaining steady improvement and reaching 85.8% accuracy. At 200B tokens, MEAP approaches optimal performance, attaining 98.2% accuracy, while the NTP baseline still falls short of 90% accuracy. It is noteworthy that MEAP achieves 85.8% accuracy using just 60B training tokens, whereas NTP requires approximately three times as many (200B tokens) to reach a similar level. This demonstrates MEAP’s superior data efficiency over NTP in key information retrieval.
@@ 4.1.3 Multi-Document Question Answering
We evaluate the model’s ability to retrieve information from long contexts using a multi-document QA task (Liu et al., 2024b) based on NaturalQuestions-Open (Kwiatkowski et al., 2019). The task requires identifying answers from a target document while ignoring $k-1$ distractor documents, where $k$ is the total number of documents in the context. We analyze performance across two context lengths with $k=10$ and $k=20$ documents and multiple answer positions. For each query, we construct an input context containing one target document with the annotated answer and $k-1$ distractor documents that do not contain any of the answers. 64K-extended models are used for this evaluation.
@@ 4.1.5 Contextual Hallucination Evaluation
Table 4: Accuracy (i.e., free of hallucinations) on text summarization datasets evaluated by different LLM judges.
@@
NTP (Deepseek-V3) 0.09 0.17 0.24
MEAP (Deepseek-V3) 0.13 0.19 0.33
NTP (Qwen-Plus) 0.16 0.11 0.21
MEAP (Qwen-Plus) 0.19 0.14 0.27
NTP(GPT-4o) 0.14 0.10 0.19
MEAP(GPT-4o) 0.16 0.13 0.24
Since MEAP improves more accurate key information retrieval, we expect it to suffer less from contextual hallucination. To verify, we evaluate MEAP in reducing contextual hallucinations on three summarization datasets: XSum (Narayan et al., 1808), WikiSum (Cohen et al., 2021), and MultiNews (Fabbri et al., 2019), following Ye et al. (2024). For this setting, we fine-tune the pre-trained models with Alpaca and evaluate them. We compare model-generated and reference summaries using (Deepseek-V3 (Liu et al., 2024a), Qwen-Plus (Yang et al., 2024) and GPT-4o (Hurst et al., 2024)) as the hallucination detector across 100 random samples per dataset. As shown in Table 4, our masking strategy achieves a consistent reduction in hallucination rates across all datasets.
@@ 4.2.1 Language Modeling Evaluation
4.2.2 Cross-Model Generalizability
To verify that MEAP’s effectiveness generalizes across different model architectures and scales, we conducted experiments on a diverse set of pre-trained LLMs. Table 6 presents results across various commonsense reasoning tasks, while Table 7 shows performance on multi-document QA tasks.
Table 6: Cross-Model Generalizability. Fine-tuning performance comparison of MEAP and NTP on commonsense reasoning tasks when applied to different architectures and model sizes.
Model Method ARC-c ARC-e BoolQ PIQA HellaSwag WinoGrande OBQA Average
Llama-3.2-3B NTP 47.95 69.07 75.54 76.50 72.43 64.33 44.40 64.32
Llama-3.2-3B MEAP 49.32 73.06 71.80 77.53 74.26 68.51 44.60 65.58
Qwen2.5-14B NTP 53.67 74.71 86.73 77.64 78.44 68.19 48.00 69.63
Qwen2.5-14B MEAP 56.83 79.38 87.37 79.33 79.37 72.69 47.40 71.77
Mistral-7B-0.2 NTP 35.67 60.10 75.81 71.22 63.03 61.40 35.40 57.52
Mistral-7B-0.2 MEAP 37.20 59.18 72.63 73.50 64.08 61.17 35.60 57.62
The results demonstrate that MEAP consistently matches or outperforms NTP across different model architectures and sizes. On the multi-document QA task, MEAP demonstrates substantial improvements across all model architectures and sizes. These results highlight MEAP’s universal effectiveness in enhancing information retrieval capabilities.
Table 7: Cross-Model Generalizability. Accuracy (%) of MEAP and NTP on multi-document QA with 20 documents across different model architectures and sizes.
Model Method Position 1 Position 5 Position 10 Position 15 Position 20 Average
Llama-3.2-3B NTP 13.60 12.09 12.54 12.69 14.35 13.05
Llama-3.2-3B MEAP 23.47 20.34 20.38 21.96 23.65 21.96
Mistral-7B-0.2 NTP 36.96 30.55 27.82 27.55 38.79 32.33
Mistral-7B-0.2 MEAP 37.91 32.98 31.46 32.22 43.45 35.60
Qwen2.5-14B NTP 60.00 51.98 56.01 56.05 63.39 57.49
Qwen2.5-14B MEAP 61.69 53.71 57.21 56.65 66.29 59.11
@@ 5 Why Does MEAP Work?
While effective, attention mechanisms in LLMs often struggle with long-context understanding, where redundant and non-informative attention is assigned to tokens (Liu et al., 2024b; Li et al., 2024b). A plausible explanation is that the attention module relies on the Softmax function to normalize attention scores within (0, 1), which tends to minimize differences among tokens, especially when training on sequences of thousands of tokens. This bears some resemblance to Martins et al. (2020)’s findings on sparse attention mechanisms which implicitly relate to their core mechanism (reducing attention to tokens).
Furthermore, LLMs exhibit a phenomenon known as attention sinks, where the initial few tokens receive disproportionately high attention scores compared to the rest (Xiao et al., 2023). Collectively, these factors lead to small and nearly indistinguishable attention scores across tokens, which is generally undesirable. When LLMs fail to properly differentiate between tokens, they are more likely to generate incorrect outputs.
@@ 5.2 MEAP Focus More on Task-Relevant Tokens
Table 10: Statistical analysis of attention score patterns between NTP and MEAP.
Length Metric Value T-Stat/P-Value
1024 Score Decay 34.08% -25.71/$<$1e-6
1024 Var. Increase 12.66% 12.26/$<$1e-6
4096 Score Decay 53.34% -9.97/$<$1e-6
4096 Var. Increase 7.80% 5.22/$<$1e-6
@@ 6 Ablation Study
6.1 Effect of Different Masking Strategies
To further investigate the impact of masking patterns, we compare three distinct masking strategies: Random Masking (our default approach), 5-Span Masking (consecutive spans of 5 tokens), and 50-Span Masking (longer spans of 50 consecutive tokens). We evaluate these strategies using a 0.3B parameter model pre-trained on 5B tokens, with results presented in Table 15.
@@ A.4 Pretrained Model Evaluation Under Different Masking Rates
Table 15: Performance comparison of different masking strategies against NTP baseline.
Method ARC-c ARC-e BoolQ PIQA HellaSwag WinoGrande OBQA Average MDQA
NTP-0.3B 18.00 37.75 58.44 62.62 28.56 50.67 13.60 40.09 0.187
Random Mask 21.84 35.44 61.25 61.04 29.50 51.46 27.40 41.13 0.218
5-Span Mask 21.42 35.61 60.40 62.08 29.81 51.07 27.60 41.14 0.168
50-Span Mask 23.46 36.20 59.54 62.84 30.43 50.99 28.00 41.64 0.189